OpenStudy (anonymous):

-2y - 7 > 11 Part 1: Solve the inequality above. Part 2: Describe the graph of the solution.

OpenStudy (michele_laino):

please add 7 to both sides of your inequality

OpenStudy (anonymous):

y>-9

OpenStudy (michele_laino):

@hislightningbug812 are you sure?

OpenStudy (anonymous):

5y > 18?

OpenStudy (michele_laino):

no, sorry @RedBlade127 -2y-7+7>11+7 so?

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

you add 7 to both sides then divide both sides by -2

OpenStudy (anonymous):

-2y - 14 > 18?

OpenStudy (michele_laino):

please note that -7+7=0, so?

OpenStudy (anonymous):

damn

OpenStudy (anonymous):

-2y-7 would be -2y

OpenStudy (michele_laino):

@RedBlade127 please retry!

OpenStudy (anonymous):

-2y >18?

OpenStudy (anonymous):

y>-9 as i said before

OpenStudy (michele_laino):

that's right!, now divide both sides by 2

OpenStudy (anonymous):

now divide both sides by -2

OpenStudy (anonymous):

no u need to divide by -2

OpenStudy (anonymous):

-2y - 7 > 11 implies -2y>11+7 implies 2y<-(11+7) implies y<-9

OpenStudy (michele_laino):

@RedBlade127 please write your answer

OpenStudy (anonymous):

uhh, I'm so confused...

OpenStudy (michele_laino):

please you have to perform this calculus: \[\frac{ -2y }{ 2 }>\frac{ 18 }{ 2 }\]

OpenStudy (anonymous):

what are your options

OpenStudy (anonymous):

@RedBlade127 remember a golden rule, whenever you multiply or divide an inequality by negative number, the sign of the inequality changes

OpenStudy (anonymous):

answer is y<-9

OpenStudy (anonymous):

thats true i forgot so y<-9 thats your answer

OpenStudy (anonymous):

Ohh, Thanks Prince and everyone.

OpenStudy (michele_laino):

@RedBlade127 thanks!